Two blocks of weights W1 = W2 = 10KN rest on a rough inclined plane and are connected by a short piece of string as shown. If the coefficient of friction is μ1 = 0.20 and μ2 = 0.30 respectively: Determine a) The inclination of the plane for which sliding will impend.



1
Expert's answer
2021-04-16T07:24:02-0400


Sliding will start when the inclination will exceed some angle "\theta".

Write forces acting on each block:


"1.\space x:f_1+T-W\sin\theta=0,\\
y: W\cos\theta=N.\\\space\\
2.\space x:f_2-T-W\sin\theta=0,\\
y: W\cos\theta=N."

Since "f=\mu N",


"1.\space \mu_1W\cos\theta+T-W\sin\theta=0,\\
2.\space \mu_2W\cos\theta-T-W\sin\theta=0,\\"

add these equations:


"\cos\theta ·W(\mu_1+\mu_2)-2W\sin\theta=0,\\
\cos\theta ·(\mu_1+\mu_2)=2\sin\theta,\\\space\\
\tan\theta=\frac{\mu_1+\mu_2}{2},\\\space\\
\theta=\arctan\frac{\mu_1+\mu_2}{2}=14°."